OpenStudy (anonymous):

f (x) = 2x^3 3xlog(x) Prove f(x) is O(x^3) by finding values for C and k. I am stuck trying to find the appropriate value for k and C.

OpenStudy (anonymous):

not sure what \(C\) and \(k\) means but if i recall this means that \[\lim_{x\to \infty}\frac{2x^3+3x\ln(x)}{x^3}=C<\infty\]